OpenStudy (anonymous):

Which statement is true about the solution if the equation 3y + 4 = -2? A. It has no solution B. It had one solution C. It has two solution D. It has infinitely many solutions

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

Here's a way of thinking about the question. When you graph a "solution", you get where the graph(s) intersect, right? This is because the points of intersection is where the (x, y) satisfy the equal sign. When you had two non-parallel lines, you had one solution because they intersect at one point. However, if you only have one line, like here, and nothing else, all you have to do is find (x, y) combinations that satisfy the equation.

OpenStudy (anonymous):

Does that in any way help you figure out the question?

OpenStudy (anonymous):

nooo :/

OpenStudy (anonymous):

Wait, sorry, I misread the question, I think. What level math are you in?

OpenStudy (anonymous):

9th

OpenStudy (anonymous):

Ok, ignore my explanation up there then. You're trying to find y-values that satisfy the equation 3y + 4 = -2. Do you remember how to solve an equation like this for a variable?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Ok, basically, you're trying to make the left side equal the right side. So, 3y + 4 has to equal -2, and you're trying to solve for what 'y' is to make this happen. The way you go about solving this is to do the opposite of order of operations. \[3y+4=-2\]\[3y+4\color{red}{-4}=-2\color{red}{-4}\]\[3y=-6\]\[3y\color{red}{\div3}=-6\color{red}{\div3}\]\[y=-2\]Now, we can check to see if this answer makes sense by plugging it back in for y. \[3(-2)+4=-6+4=-2\]Because we got -2, this y-value is correct, and it is the only y-value that works. If you look, we only got 1 y-value, and the number of y-values is the number of solutions you have.

OpenStudy (anonymous):

so one solution ?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

thank youu